OpenStudy (anonymous):

Factor the trinomial x2 - x - 6. Write each factor as a polynomial in descending order.

OpenStudy (campbell_st):

well it has 2 factors... find the factors of -6 that add to -1 the larger factor is negative... the smaller positive then write it as (x + factor 1)(x + factor 2) hope it helps

OpenStudy (campbell_st):

there you go... always someone will provide the answer

OpenStudy (anonymous):

x2 - x - 6 =x2 -3x +2x -6 =x(x-3) +2(x-3) =(x+2) (x-3)